OpenStudy (anonymous):

please help?! simplify the expression. (-2i)(8i)

OpenStudy (solomonzelman):

\[(-2i)(8i)=(-2) \times 8 \times i \times i=(-16) \times i^2= (-16) \times (-1)=16\]

OpenStudy (anonymous):

i^2 = -1

OpenStudy (anonymous):

wow... I really always overthink these problems. I had a feeling that was all I had to do. thank you, I will remember to simplify this way.

OpenStudy (anonymous):

yea man it's easier than you think bro

OpenStudy (anonymous):

looks like it! thanks guys